. A 50 kg mass is sitting on a frictionless surface. An unknown constant force pushes the mass for


2 seconds until the mass reaches a velocity of 3 m/s


a. What is the initial momentum of the mass?


b. What is the final momentum of the mass?


c. What was the force acting on the mass?


d. What was the impulse acting on the mass?

1
Expert's answer
2022-02-11T08:46:10-0500

Given:

"m=50\:\rm kg"

"v_1=0"

"v_2=3\:\rm m/s"

"t=2\:\rm s"


a. The initial momentum of the mass

"p_1=mv_1=0"

b. The final momentum of the mass

"p_2=mv_2=\rm 50*3=150\: kg*m/s"

c. The force acting on the mass

"F=\frac{\Delta p}{\Delta t}=\frac{150-0}{2}=75\:\rm N"

d. The impulse acting on the mass

"\Delta p=150\:\rm kg*m/s"